# Learning with 3D rotations, a hitchhiker’s guide to SO(3)

### v1

@@ 2 Representations of rotation
Figure 3: Exponential coordinates (Exp) and axis-angle representation and their double cover property. Top: EC: rotation around $\omega$ by angle $\|\omega\|$. The vector $\omega_{1}=\alpha_{1}\tilde{\omega}_{1}\in\mathbb{R}^{3}$ describes the same rotation as $\omega_{2}=(\|\omega_{1}\|-2\pi)\omega_{1}/\|\omega_{1}\|$. Bottom: Axis-angles explicitly represent axis and angle: $\tilde{\omega}\in\mathcal{S}^{2}$, $\alpha\in\mathbb{R}$. The vector $[\tilde{\omega},\alpha]$ describes the same rotation as $[-\tilde{\omega},-\alpha]$.
@@ 4.1 Why does rotation estimation need high-dimensional rotation parameterization?
Distance-picking and measuring distances in $\mathrm{SO}(3)$ are common strategies found in literature related to rotation estimation. In distance picking, we choose the minimum distance between elements in $\mathcal{R}$ taking double cover into account e.g. Equations 5 and 6. Alternatively, we can directly compute distances in $\mathrm{SO}(3)$, e.g. via $d_{c}(f(r_{1}),f(r_{2}))$, such that the gradients $\nabla_{y}d_{c}(f(y),f(r_{2}))$ point in directions in $\mathcal{R}$ that increase the metric in $\mathrm{SO}(3)$. However, from the previous discussion and the additional argument provided in Section C.5, we conclude the following key insight:
@@
Brégier (2021) defined the loss function $L=v_{1}^{T}f(r)v_{2}$ with $v_{i}$ being vectors pointing in uniform random direction and $r$ denoting either quaternions, $\mathbb{R}^{6}$+GSO, or $\mathbb{R}^{9}$+SVD. For random representation vectors $x\sim\mathcal{N}(0,1)^{n}$, this loss function has a smaller absolute error to its linear approximation using $\mathbb{R}^{9}$+SVD compared to using $\mathbb{R}^{6}$+GSO or quaternions. Inspired by Brégier (2021), (Dinh et al., 2017), (Gilmer et al., 2021), and Lyle et al. (2023), we take a closer look at the gradients of the loss function
@@
with the target rotation $R:=I_{3}$ and $f(r)$ denoting either an $\mathbb{R}^{6}$+GSO or $\mathbb{R}^{9}$+SVD layer. Assume a network predicts $r$ such that its parameters $\theta$ could be updated via the gradient $\nabla_{r}L\circ\nabla_{\theta}r$. Then, we are interested in how $\nabla_{r}L$ varies when using either $\mathbb{R}^{6}$+GSO or $\mathbb{R}^{9}$+SVD. To illustrate this, we plot in Figure 10 the path of the representation vectors of $\mathbb{R}^{6}$+GSO and $\mathbb{R}^{9}$+SVD when optimizing (15) using gradient-descent with momentum. The interested reader is pointed to the python script ‘gso_vs_svd_animation” that created Figure 10. Due to the way GSO and SVD construct $f(r)$, we observed instabilities in the above optimization for $\nu_{1}\approx e_{3}$, $\nu_{1}||\nu_{2}$, $\text{det}(M)=0$, and $M=-R$.
Moreover, the speed of convergence is notably affected by large ratios between the length of the representation vectors which particularly seems to affect $\mathbb{R}^{6}$+GSO. We compute the ratios between the length of the gradients of (15) wrt. $\nu_{1},\nu_{2}$ or $m_{1},m_{2},m_{3}$. As illustrated in Figure 11 for 20000 uniform randomly initialized $r\in[-2,2]$, the gradient ratios remain notably closer to one for $\mathbb{R}^{9}$+SVD compared to $\mathbb{R}^{6}$+GSO. Further, in $\mathbb{R}^{9}$+SVD, the distributions of gradient ratios are similar. We conclude that in the hypercube with a side-length of four in which we sampled $r$, considerably more instances of $r$ exhibit gradients that impede training when using $\mathbb{R}^{6}$+GSO compared to $\mathbb{R}^{9}$+SVD.
@@
Another argument why SVD performs better than GSO evolves around the higher-dimensionality of the former. In various fields such as position encodings (Dufter et al., 2022), Koopman operator theory (Brunton et al., 2021), and neural ODEs (Zhang et al., 2020), it has been observed that higher dimensional representations may benefit learning. This is particularly the case if the function we discern from data resides in a three-dimensional manifold such as $\mathrm{SO}(3)$.
@@ 4.2 What to be aware of for feature prediction?
Figure 12: For $\mathrm{SO}(2)$, the function $\phi=f(\cos(\phi),\sin(\phi))$ is continuous such that feature prediction is feasible. Albeit, connected sets in $\mathrm{SO}(2)$ may not be connected in $\mathcal{R}$ (red lines). The effect of disconnected pre-images can be reduced through data augmentation (orange lines).
@@
Again, they work out of the box. They also perform better in practice, as we show in Section 5, because more input dimensions allow for a simpler functional form for $h$. A relation can be drawn to position encodings used in transformers (Dufter et al., 2022). A sin/cos representation (and potentially higher frequencies) works better than using a scalar, although it contains the same information in principle.
@@ 5 Experiments
The task is to predict from two point clouds $P_{1},P_{2}\in\mathbb{R}^{N}$, with $N=3000$, the rotation $R$ such that $RP_{2}=P_{1}$ (Zhou et al., 2019). We follow the same model and dataset generation pipeline as Levinson et al. (2020). Point clouds are extracted from a set of $726$ airplane CAD models and undergo uniformly sampled rotations in $SO(3)$ to eliminate any potential rotation bias in the dataset. For testing, 100 pairs are left out. Figure 13 shows that $\mathbb{R}^{9}$+SVD is best followed by $\mathbb{R}^{6}$+GSO. Figure 23 shows that for quaternions deploying a half-space map notably improves performance whereas distance picking reduces it.
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Figure 14: Experiment 2. Left: Rotation estimation from cube image where $\mathbb{R}^{9}$+SVD performs best. See Figure 24 for additional results. Right: Network trained with pixel-MSE loss renders a cube from a given rotation. While quaternion augmentation slightly reduces the error, $\mathbb{R}^{9}$+SVD clearly outperforms the other representations (note the logarithmic axis scale).
@@
In this experiment, we predict the orientation of a colorful cube represented as a rendered image. The experimental setup is explained in Section D.2. The results in Figure 14 (left) paint the the same picture as before showing that $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO perform notably better than low-dimensional representations.
@@
Now, we look at the inverse problem to Experiment 2.1 predicting the cube’s image from an representation of its orientation. The experiment is further described in Section D.2. The results in Figure 14 (right) show that $\mathbb{R}^{9}$+SVD performed notably better than $\mathbb{R}^{6}$+GSO while quaternions benefit from deploying a half-space map.
@@
In this experiment, we re-train the network of Wang et al. (2019) using different rotation representations. Simply put, the network predicts a translation and rotation representation which is used to transform an object’s point cloud such that it aligns with the ground truth point cloud. This model has become a common baseline for object pose estimation from a single RGB-D camera image. Figure 15 shows the average accuracy across all objects for three seeds. Euler angles perform notably worse than $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO whereas quaternions perform similarly. We hypothesise, that quaternions perform well due to the data set mostly containing small angles. For details, see Section D.3.
[…]
@@
with period $L=2$, fourier parameters $A_{k},B_{k}$, and $t:\mathrm{SO}(3)\to\mathbb{R}$ being a 2-layer MLP with ReLu activations. Both the parameters of $t(\vec{(}R))$ and the Fourier parameters are randomly initialized. If we increase $n_{b}$, then the complexity of (16) increases in expectation. As further detailed in Sec. D.4, we approximate each randomly generated target function with an MLP which has a rotation representation as input. Figure 16 shows the training results from $n_{b}=1$ up to $n_{b}=7$ for 100 target functions each. We keep the number of points fixed to 800, 400, and 1000 for the train, the validation, and the test data set, respectively. As we increase $n_{b}$, the target function becomes more wiggly on average and the model MSEs increase. Augmentation of quaternions (augmentation threshold b=0.1) has a notable effect for $n_{b}=1$ to $n_{b}=3$. Albeit, for larger $n_{b}$, the error arising from having no data points beyond the domain boundaries seem to be of minor importance compared to other errors affecting the canonical quaternions ($\mathrm{Q}^{\mathrm{+}}$) such that they compare similarly to augmented quaternions.
@@ Appendix A Notation
Vectors are denoted as $a=[a_{1},a_{2},a_{3}]\in\mathbb{R}^{3\times 1}$. The identity matrix of size three is denoted as $I_{3}$. We denote the $n$ dimensional unit sphere as $\mathcal{S}^{n}=\{x\in\mathbb{R}^{n+1}:\|x\|=1\}$. Given a vector $a\in\mathbb{R}^{d}$, $\mathrm{diag}(a)\in\mathbb{R}^{d\times d}$ denotes a diagonal matrix with the entries of $a$ on its diagonal. To create a skew-symmetric matrix from a vector $a=[a_{1},a_{2},a_{3}]\in\mathbb{R}^{3}$, we have
@@ Appendix B Matrix representation of SO(3)
[$\ m_{1}&m_{2}&m_{3}\\ \mid&\mid&\mid\end{bmatrix}\in\mathbb{R}^{3\times 3}$ (18)
@@
As illustrated in Fig. 17, rotations are commonly represented by a vector $a\in\mathbb{R}^{3}$ transformed by matrix $M$ to produce $b=Ma\in\mathbb{R}^{3}$ where $\mathbb{R}^{3}$. Alternatively, we gain geometric intuition on the rotation matrix itself if we imagine the vector $a$ to be described wrt. to a frame $\{A\}$, writing ${}^{A}a=[a_{1},a_{2},a_{3}]$, and the vector ${}^{B}a$ being obtained as
@@
Appendix C Additional aspects on learning with rotations
C.1 Intuition on some terminology from linear algebra
@@
“Special” refers to $\det(M)=1$. If $\det(M)=1$, then the column vectors of $M$ form a right-hand coordinate system. Orthogonal matrices with determinant of one are called proper rotations, whereas orthogonal matrices with determinant of negative one are called improper rotations. Improper rotations combine rotation with reflection and do not form a group as the product between two improper rotations is a proper rotation. That said, we encounter improper rotations in the description of rotations via geometric algebra where rotations are described by means of two reflections.
C.2 Why learn with rotation representations that have three or four parameters?
@@
C.3 How do we augment quaternions for feature prediction?
@@
C.4 Are rotors underrated?
Pepe et al. (2022) contends that using bi-vectors (3D) and rotors (4D) for rotation estimation might rival rotation representations employing the Gram-Schmidt orthonormalization (6D). However, bi-vectors / rotors are in fact computationally equivalent to exponential coordinates / quaternions as shown in a detailed code comparison by Bosch (2020a). While a geometric algebra perspective on rotations (including rotors and bi-vectors) may be more intuitive than using quaternions, as exemplified by Macdonald (2010, p. 87-91) and Bosch (2020b), we are skeptical about significant performance differences for rotation estimation.
@@ C.4 Are rotors underrated?
We advocate for further analysis that considers the proximity of representation vectors to singularities, measures all errors in $\mathrm{SO}(3)$ rather than in rotor space, and tries using a half-space map for quaternions and exponential coordinates. One experiment found in Zhou et al. (2019); Levinson et al. (2020); Brégier (2021); Pepe et al. (2022) is “Inverse Kinematics with CMU MoCap data” (see Sec. D.5), for which we analysed how far the rotations are from the unit rotation. We hypothesized that for some joints the anatomy of human kinematics does not support large angle ranges and in turn the data set contains mostly small rotations (recall that a rotation $R\in\mathrm{SO}(3)$ is small if $\|\mathbb{I}-R\|_{\mathrm{F}}\leq\sqrt{2}$). Indeed, as shown in Figure 19, this data set contains significantly more small rotations that are close to the unit rotation which reduce the effect of singularities on learning with rotation representations with three or four dimensions.
@@
C.5 How does distance picking and computing distances in $\mathrm{SO}(3)$ affect rotation estimation?

### v2

@@ 2 Representations of rotation
Figure 3: Exponential coordinates (Exp) and axis-angle representation and their double cover property. Top: Exp. coord.: rotation around $\omega$ by angle $\|\omega\|$. The vector $\omega_{1}=\alpha_{1}\tilde{\omega}_{1}\in\mathbb{R}^{3}$ describes the same rotation as $\omega_{2}=(\|\omega_{1}\|-2\pi)\omega_{1}/\|\omega_{1}\|$. Bottom: Axis-angles explicitly represent axis and angle: $\tilde{\omega}\in\mathcal{S}^{2}$, $\alpha\in\mathbb{R}$. The vector $[\tilde{\omega},\alpha]$ describes the same rotation as $[-\tilde{\omega},-\alpha]$.
@@ 4.1 Why does rotation estimation need high-dimensional rotation parameterization?
Distance-picking and measuring distances in $\mathrm{SO}(3)$ are common strategies found in literature related to rotation estimation. In distance picking, we choose the minimum distance between elements in $\mathcal{R}$ taking double cover into account e.g. Equations 5 and 6. Alternatively, we can directly compute distances in $\mathrm{SO}(3)$, e.g. via $d_{c}(f(r_{1}),f(r_{2}))$, such that the gradients $\nabla_{y}d_{c}(f(y),f(r_{2}))$ point in directions in $\mathcal{R}$ that increase the metric in $\mathrm{SO}(3)$. However, from the previous discussion and the additional argument provided in Section D.4, we conclude the following key insight:
@@
Brégier (2021) defined the loss function $L=v_{1}^{T}f(r)v_{2}$ with $v_{i}$ being vectors pointing in uniform random direction and $r$ denoting either quaternions, $\mathbb{R}^{6}$+GSO, or $\mathbb{R}^{9}$+SVD. For random representation vectors $r\sim\mathcal{N}(0,1)^{n}$, this loss function has a smaller absolute error to its linear approximation using $\mathbb{R}^{9}$+SVD compared to using $\mathbb{R}^{6}$+GSO or quaternions. Inspired by Brégier (2021), (Dinh et al., 2017), (Gilmer et al., 2022), and Lyle et al. (2023), we take a closer look at the gradients of the loss function
@@
with the target rotation $R:=\mathbb{I}$ and $f(r)$ denoting either an $\mathbb{R}^{6}$+GSO or $\mathbb{R}^{9}$+SVD layer. Assume a network predicts $r$ such that its parameters $\theta$ could be updated via the gradient $\nabla_{\theta}r\circ\nabla_{r}L$. Then, we are interested in how $\nabla_{r}L$ varies when using either $\mathbb{R}^{6}$+GSO or $\mathbb{R}^{9}$+SVD. To illustrate this, we plot in Figure 10 the path of the representation vectors of $\mathbb{R}^{6}$+GSO and $\mathbb{R}^{9}$+SVD when optimizing (15) via gradient-descent with momentum. Due to the way GSO and SVD construct $f(r)$, we observed instabilities in the optimization for $\nu_{1}\approx e_{3}$, $\nu_{1}||\nu_{2}$, $\text{det}(M)=0$, and $M=-R$.
Moreover, the speed of convergence is notably affected by large ratios between the length of the representation vectors which particularly seems to affect $\mathbb{R}^{6}$+GSO. We compute the ratios between the length of the gradients of (15) wrt. $\nu_{1},\nu_{2}$ or $m_{1},m_{2},m_{3}$. As illustrated in Figure 11 for 20000 uniform randomly initialized $r\in[-2,2]$, the gradient ratios remain notably closer to one for $\mathbb{R}^{9}$+SVD compared to $\mathbb{R}^{6}$+GSO. Further, in $\mathbb{R}^{9}$+SVD, the distributions of gradient ratios are similar. We conclude that in the hypercube with a side-length of four in which we sampled $r$, when using $\mathbb{R}^{6}$+GSO compared to $\mathbb{R}^{9}$+SVD considerably more instances of $r$ exhibit gradients that impede training.
@@
Another argument why SVD performs better than GSO evolves around the higher-dimensionality of the former. In various fields such as position encodings (Dufter et al., 2022), Koopman operator theory (Brunton et al., 2022), and neural ODEs (Zhang et al., 2020), it has been observed that higher dimensional representations may benefit learning. This is particularly the case if the function we discern from data resides in a three-dimensional manifold such as $\mathrm{SO}(3)$.
@@ 4.2 What to be aware of for feature prediction?
Figure 12: For $\mathrm{SO}(2)$, the function $\phi=f(\cos(\phi),\sin(\phi))$ is continuous, rendering feature prediction feasible. However, some connected sets in $\mathrm{SO}(2)$ are not connected in $\mathcal{R}$ (red). Data augmentation (orange) mitigates the effect of disconnected pre-images.
@@
Again, they work out of the box and perform better in practice, as we show in Section 5. More input dimensions allow for a simpler functional form for $h$. A relation can be drawn to position encodings used in transformers (Dufter et al., 2022). A sin/cos representation (including higher frequencies) works better than using a scalar.
@@ 5 Experiments
The task is to predict from two point clouds $P_{1},P_{2}\in\mathbb{R}^{N}$, with $N=3000$, the rotation $R$ such that $RP_{2}=P_{1}$ (Zhou et al., 2019). We follow the model and dataset generation pipeline as Levinson et al. (2020). Point clouds are extracted from a set of $726$ airplane CAD models and undergo uniformly sampled rotations in $SO(3)$ to eliminate any potential rotation bias in the dataset. For testing, 100 pairs are left out. Figure 13 shows that $\mathbb{R}^{9}$+SVD is best followed by $\mathbb{R}^{6}$+GSO. Figure 23 shows that for quaternions deploying a half-space map notably improves performance whereas distance picking reduces it.
@@
Figure 14: Experiment 2. Left: Rotation estimation from cube image where $\mathbb{R}^{9}$+SVD performs best. See Figure 24 for additional results. Right: Network trained with pixel-MSE loss renders a cube from a given rotation. While quaternion augmentation slightly reduces the error, $\mathbb{R}^{9}$+SVD shows superior performance.
@@
In this experiment, we predict the orientation of a colorful cube represented as a rendered image. The experimental setup is explained in Section E.2. The results in Figure 14 (left) paint the the same picture as before showing that $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO perform notably better than low-dimensional representations.
@@
Now, we look at the inverse problem to Experiment 2.1 predicting the cube’s image from an representation of its orientation. The experiment is further described in Section E.2. The results in Figure 14 (right) show that $\mathbb{R}^{9}$+SVD performed notably better than $\mathbb{R}^{6}$+GSO while quaternions benefit from deploying a half-space map.
@@
In this experiment, we re-train the network of Wang et al. (2019) using different rotation representations. Simply put, the network predicts a translation and rotation representation which is used to transform an object’s point cloud such that it aligns with the ground truth point cloud. This model has become a common baseline for object pose estimation from a single RGB-D camera image. Figure 15 shows the average accuracy across all objects for three seeds. Euler angles perform notably worse than $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO whereas quaternions perform similarly. We hypothesise, that quaternions perform well due to the data set mostly containing small angles. For details, see Section E.3.
@@
with period $L=2$, fourier parameters $A_{k},B_{k}$, and $t:\mathrm{SO}(3)\to\mathbb{R}$ being a 2-layer MLP with ReLu activations. Both the parameters of $t(\mathrm{vec}(R))$ and the Fourier parameters are randomly initialized. If we increase $n_{b}$, then the complexity of (16) increases in expectation. As further detailed in Sec. E.4, we approximate each randomly generated target function with an MLP which has a rotation representation as input. Figure 16 shows the training results from $n_{b}=1$ up to $n_{b}=5$ for 100 target functions each. We keep the number of points fixed to 800, 200, and 1000 for the train, the validation, and the test data set, respectively. As we increase $n_{b}$, the target function becomes on average more wiggly and the model MSEs increase. Augmentation of quaternions (augmentation threshold b=0.1) has a notable effect for $n_{b}=1$ to $n_{b}=3$. Albeit, for larger $n_{b}$, the error arising from having no data points beyond the domain boundaries seem to be of minor importance.
[…]
@@ 6 Conclusion
Numerous extensions to the discussed topics are being proposed. Notably, Peretroukhin et al. (2020) represents rotations through a symmetric matrix that defines an antipodally symmetric distribution over quaternions effectively mitigating the problems arising due to double cover. This approach outperformed $\mathbb{R}^{6}$+GSO on various benchmarks. Chen et al. (2022) notably improved prediction performance of $\mathbb{R}^{6}$+GSO, $\mathbb{R}^{9}$+SVD, and (Peretroukhin et al., 2020) layers whose predictions are fed into an L2 loss by adjusting the gradients during the backward pass which marginally increases computation times.
@@ Acknowledgments
This work was supported by the ERC - 101045454 REAL-RL and the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039B). Georg Martius is a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645.
@@ Appendix A Notation
Vectors are denoted as $a=[a_{1},a_{2},a_{3}]\in\mathbb{R}^{3\times 1}$. The identity matrix of size three is denoted as $\mathbb{I}$. We denote the $n$ dimensional unit sphere as $\mathcal{S}^{n}=\{x\in\mathbb{R}^{n+1}:\|x\|=1\}$. Given a vector $a\in\mathbb{R}^{d}$, $\mathrm{diag}(a)\in\mathbb{R}^{d\times d}$ denotes a diagonal matrix with the entries of $a$ on its diagonal. Given a vector $a=[a_{1},a_{2},a_{3}]\in\mathbb{R}^{3}$, $a\times a=[a]_{\times}\cdot a$ where
@@ Appendix B Matrix representation of SO(3)
$M=\begin{bmatrix}\mid&\mid&\mid\\ m_{1}&m_{2}&m_{3}\\ \mid&\mid&\mid\end{bmatrix}\in\mathbb{R}^{3\times 3}$ (18)
@@
As illustrated in Fig. 17, rotations are commonly represented by a vector $a\in\mathbb{R}^{3}$ transformed by matrix $M$ to produce $b=Ma\in\mathbb{R}^{3}$. Alternatively, we gain geometric intuition on the rotation matrix itself if we imagine the vector $a$ to be described wrt. to a frame $\{A\}$, writing ${}^{A}a=[a_{1},a_{2},a_{3}]$, and the vector ${}^{B}a$ being obtained as
@@
Appendix C Intuition on some terminology from linear algebra
@@
“Special” refers to $\det(M)=1$. If $\det(M)=1$, then the column vectors of $M$ form a right-hand coordinate system. Orthogonal matrices with determinant of one are called proper rotations, whereas orthogonal matrices with determinant of negative one are called improper rotations. Improper rotations combine rotation with reflection, the product between two improper rotations is a proper rotation. We encounter improper rotations in the description of rotations via geometric algebra where rotations are described by means of two reflections.
Appendix D Additional aspects on learning with rotations
D.1 Why learn with rotation representations that have three or four parameters?
@@
D.2 How do we augment quaternions for feature prediction?
@@
D.3 Are rotors underrated?
Pepe et al. (2022) contends that using bi-vectors (3D) and rotors (4D) for rotation estimation might rival rotation representations employing the Gram-Schmidt orthonormalization (6D). However, bi-vectors / rotors are in fact computationally equivalent to exponential coordinates / quaternions as shown in a detailed code comparison by Bosch (2020a). While a geometric algebra perspective on rotations (including rotors and bi-vectors) may be more intuitive than using quaternions, as exemplified by Macdonald (2010, p. 87-91) and Bosch (2020b), we remain cautious about expecting significant performance differences for rotation estimation.
@@ D.3 Are rotors underrated?
We advocate for further analysis that considers the proximity of representation vectors to singularities, measures all errors in $\mathrm{SO}(3)$ rather than in rotor space, and tries using a half-space map for quaternions and exponential coordinates. One experiment found in Zhou et al. (2019); Levinson et al. (2020); Brégier (2021); Pepe et al. (2022) is “Inverse Kinematics with CMU MoCap data” (see Sec. E.5), for which we analysed how far the rotations are from the unit rotation. We hypothesized that for some joints the anatomy of human kinematics does not support large angle ranges and in turn the data set contains mostly small rotations (recall that a rotation $R\in\mathrm{SO}(3)$ is small if $\|\mathbb{I}-R\|_{\mathrm{F}}\leq\sqrt{2}$). Indeed, as shown in Figure 19, this data set contains significantly more small rotations that are close to the unit rotation which reduce the effect of singularities on learning with rotation representations with three or four dimensions.
@@
D.4 How does distance picking and computing distances in $\mathrm{SO}(3)$ affect rotation estimation?